# DGFusion: Depth-Guided Sensor Fusion for Robust Semantic Perception

### v2

@@ I Introduction
Robust segmentation for automated driving systems demands reliable sensor fusion under any environmental condition. Conventional semantic perception pipelines that rely on a single sensor struggle in domains such as fog, rain, or low light, where the performance of individual sensors degrades. Prior work [cafuser] has shown that conditioning cross-modal fusion on a global Condition Token (CT) that encodes the scene-wide environmental condition can effectively guide fusion by dynamically adapting sensor contributions. However, while this global approach models the overall scene condition, it overlooks the spatial variability of the effect of this condition on the sensor measurements, which largely depends on the local depth.
@@
In practice, depth is a key factor affecting the variable reliability of each sensor in each part of the scene. For instance, lidars provide accurate returns for nearby objects but degrade rapidly at longer distances, e.g. in fog or snowfall [hahner2021fog, hahner2022lidar]. Radars, on the other hand, operate well at long distances even in adverse weather but suffer from strong noise due to multi-path effects, whereas cameras are reliable only in well-lit parts of the scene, which are in turn related to the 3D structure of the scene. In other words, environmental conditions introduce spatial variations in the signal-to-noise ratio of each sensor, which call for a fine-grained local adaptivity of the fusion strategy. Consequently, incorporating explicit depth cues into the fusion mechanism for semantic perception could enable such spatial adaptivity to sensor reliability, which would in turn enhance segmentation robustness in challenging environments. This hypothesis is reinforced by several works on RGB-only semantics and depth multi-tasking [bruggemann2021exploring, rahman2024semi], which improve segmentation performance over semantics-only learning. Despite this promise for depth-informed fusion, most state-of-the-art multi-sensor segmentation methods [broedermann2023hrfuser, zhang2023delivering, li2024stitchfusion, jia2024geminifusion] treat lidar solely as an input modality, overlooking its potential for improving segmentation via multi-tasking by leveraging its pixel-aligned sparse depth measurements as ground truth for learning depth together with semantics.
@@
We fill in this gap and propose DGFusion, a depth-guided multimodal fusion network which upgrades condition-aware sensor fusion for semantic perception with explicit local depth guidance (cf. Fig. 1). Building on CAFuser [cafuser], we adopt its shared backbone with modality-specific feature adapters and its global CT that encodes scene-wide environmental conditions. Beyond this, in DGFusion we introduce an auxiliary depth prediction head that leverages lidar’s pixel-aligned sparse depth measurements as noisy supervision to enrich local features with implicit depth cues to guide each local-window cross-attention fusion. This dual encoding of global environmental conditions and spatially-varying depth cues enables the network to make more precise decisions based on which sensor to rely more on for each feature. By incorporating depth cues, the fusion mechanism can adaptively adjust the weighting of each modality per image region, favoring the per-case more reliable sensor. Moreover, since these noisy depth measurements are already available as one of our network’s inputs, i.e. the lidar, depth supervision comes at no extra cost compared to existing fusion pipelines. The synergy between the global condition and local depth representation is especially beneficial in adverse conditions, which have been found [seeing_through_fog] to cause pronounced spatial variations in sensor reliability.
@@
Besides, training our model on adverse conditions poses additional challenges. For example, in fog or snowfall the lidar measurements, which provide pixel-aligned depth supervision, become inherently noisy [hahner2021fog, hahner2022lidar], making the direct integration of standard depth estimation approaches and losses in our multi-tasking model infeasible. In order to get a reliable training signal across all conditions, we carefully design our depth loss to be robust to sparse and noisy ground truth by combining an outlier-robust $L_{1}$ loss with strategies commonly used in monocular depth estimation. These strategies consist of an edge-aware smoothness loss to penalize depth gradients in uniform regions of the RGB image and a novel panoptic-edge-aware smoothness loss that better captures depth discontinuities between instances of the same semantic classes compared to previous semantic-edge-aware smoothness losses.
@@
Casting the training of DGFusion as a multi-tasking problem leverages the synergy between depth and semantics with minimal changes to the inference pipeline. Extensive experiments on MUSES [MUSES] and DeLiVER [zhang2023delivering] demonstrate that DGFusion sets the new state of the art in multi-sensor semantic perception across diverse conditions.
@@ II Related Work
Semantic perception encompasses tasks like semantic [Cityscapes], panoptic [kirillov2019panoptic], and instance [he2017mask] segmentation, which are pivotal for robotics applications such as automated driving [does:vision:matter:for:action]. Recent mask-based networks have advanced the state of the art [li2023mask], and unifying these tasks within a single model has proven beneficial [jain2023oneformer]. In contrast to these RGB-only methods, our work focuses on multimodal semantic perception, leveraging sensor fusion to enhance robustness in challenging conditions where RGB data is unreliable.
@@
Sensor fusion is central in practical vision applications, as different sensors feature complementary characteristics. The availability of large multimodal datasets, from early ones like KITTI [kitti] to recent ones focusing on adverse conditions [zhang2023delivering, seeing_through_fog, MUSES], has driven the development of fusion methods. Architectures have evolved from two-modality systems [chaturvedi_pay] to modular, attention-based networks that handle an arbitrary number of inputs [broedermann2023hrfuser, zhang2023delivering]. Recent approaches have integrated large-scale pre-trained models [li2024stitchfusion], explored modality-agnostic designs with shared backbones [centeringthevalue], and combined intra-modal with inter-modal attention [jia2024geminifusion]. The MUSES method [MUSES] extends [cheng2022masked] to the multimodal setting, solving panoptic and semantic segmentation via local-window cross-attention. However, all these works fuse sensors uniformly, lacking explicit adaptation to the conditions that affect each image region non-uniformly, largely depending on depth.
@@
Condition-aware fusion utilizes information about the environmental conditions to guide perception. [chaturvedi_pay] addresses object detection by implicitly assigning larger weights to the modality that yields better detection features at late fusion. Knowledge distillation from vision-language models is used in [learningmodalityagnostic] for learning modality-agnostic representations. The most closely related work to ours in this regard is [cafuser], which guides cross-modal attention with a global camera-based CT that is learned contrastively based on training-time supervision with image-level verbal condition descriptions. In contrast, we aim at a more fine-grained, pixel-level adaptivity to the environmental condition, as its effects on sensor data vary non-uniformly with depth.
@@
Semantics and depth are strongly correlated tasks, and it is well-established that they benefit from being learned jointly in a multi-task setup [bruggemann2021exploring, rahman2024semi]; cf. [vandenhende2021multitask] for a comprehensive overview of such semantics-depth multi-tasking methods. Prior work has leveraged this synergy, with semantics guiding monocular depth estimation [chen2019towards, zhu2020edge]. Similarly, depth supervision has benefited from edge-aware smoothness losses based on image gradients [godard2017unsupervised] and semantics-aware losses that use semantic annotations [li2024semantic], which we extend to use more fine-grained panoptic annotations. In contrast to previous works, we explore the use of depth in multimodal settings as a readily available input signal, e.g. from sparse lidar measurements in typical outdoor multi-sensor suites, in order to benefit semantics. While the aforementioned multi-sensor semantic perception works utilize such depth information only as input to the fusion architecture, we instead leverage it as a reference output during training, reframing sensor fusion as a multi-task problem where depth completion serves as an auxiliary task to learn informative, localized features for semantic perception. The internal depth features learned from this auxiliary task are leveraged as a spatially varying conditioning in our attentive fusion together with global condition information. This combination allows DGFusion to adaptively weigh modalities not only per-input, but also per-region, based on how depth modulates the effect of the environmental condition on each sensor.
@@ III-A Architecture
Our goal is to achieve robust multimodal segmentation in all weather conditions. We build on CAFuser [cafuser], which processes all modalities through a shared backbone with lightweight modality-specific feature adapters to align heterogeneous inputs into a common feature space, and uses global environmental condition encodings to guide the sensor fusion for the OneFormer [jain2023oneformer] segmentation head. Following CAFuser, we project the secondary modalities (lidar, radar, and event camera) as 3-channel images onto the RGB camera plane and dilate them with a kernel of size $K_{\text{sensor}}$ to compensate for sparsity. We adopt CAFuser’s shared backbone with feature adapters and its global condition token (CT). In contrast to CAFuser, we treat lidar not only as an input modality but also as sparse, noisy ground-truth depth for multi-task learning, which poses additional challenges in adverse conditions where lidar measurements degrade. With our carefully designed robust losses (cf. Sec. III-B), this enables our network to learn depth-informed features that provide additional supervision for semantics and enable locally adaptive depth-guided fusion via local depth tokens.
In Fig. 2, our complete DGFusion architecture is depicted. On the left, an arbitrary number of sensor modalities are processed by the shared backbone with modality-specific feature adapters, resulting in per-level feature maps. These features are then split into three branches: the depth estimation branch (top), the segmentation branch (center), and the condition representation branch adopted from CAFuser (bottom). The depth branch uses raw, undilated lidar ground truth to supervise depth predictions and induce depth-informed features via a lightweight fusion module, while the condition branch generates a global CT from the RGB input. Both branches provide supplementary cues to the central depth-guided fusion module.
@@
We use a Semantic FPN [kirillov2019panoptic] decoder as our auxiliary depth head, which upsamples each depth feature $\mathbf{d}_{l}$ to 1/4 resolution and aggregates them:
@@
where Upsample consists of multiple convolutions and bilinear upsampling steps. This structure fuses coarse and fine information efficiently, enabling robust depth estimation at various scales. At inference, we drop the auxiliary depth head but keep the depth features $\mathbf{d}_{l}$ for guiding fusion in the semantic branch.
Condition representation branch. We adopt CAFuser’s [cafuser] strategy to produce a global CT. We flatten the highest-level RGB features and process them with a lightweight Transformer composed of 2 encoder and 2 decoder layers:
@@
supervised during training with a verbo-visual contrastive loss [cafuser, jain2023oneformer, bruggemann2023contrastive] that leverages text prompts describing the environmental conditions in detail.
Depth-guided fusion. We use the depth features $\mathbf{d}_{l}$ to guide sensor fusion in our main semantic branch. Our Depth-Guided Fusion module utilizes parallel multi-window cross-attention fusion [broedermann2023hrfuser] to fuse in parallel each secondary modality via cross-attention with the primary camera modality within $K_{w}\times K_{w}$ local windows. This fusion design is efficient and scales well with the number of input modalities. For each level $l$ and secondary modality $m$, we partition RGB features $\mathbf{F}_{\text{rgb}}^{l}$, secondary features $\mathbf{F}_{\text{sec}}^{l,m}$, and depth features $\mathbf{d}_{l}$ into $I_{l}$ windows of size $K_{w}\times K_{w}$. For each window $i\in\{1,\ldots,I_{l}\}$, we extract a local Depth Token (DT):
@@ III-B Losses
Outlier-robust $L_{1}$ loss. Lidar returns degrade in adverse conditions, making the raw depth measurements noisy and unreliable. To address this, we compute an $L_{1}$ loss on logarithmic depth values. The $L_{1}$ loss is inherently more robust to outliers than $L_{2}$, and using the logarithmic scale better handles depth variations across orders of magnitude [rahman2024semi, piccinelli2024unidepth, piccinelli2025unidepthv2]. Let $\hat{D}_{p}$ be the predicted depth at pixel $p$ and $D_{p}$ the corresponding raw lidar reference depth. Since raw lidar depths are often severely corrupted in adverse conditions, directly supervising depth predictions using all available measurements can degrade learning (cf. Sec. IV-B). To address this, we restrict the loss computation to a more reliable subset of pixels, i.e. those pixels below the $\tau$-quantile of depth errors. We first define the absolute log-error at pixel $p$ as
@@
Edge-aware smoothness loss. To guide depth prediction in regions lacking reliable lidar supervision, we employ an edge-aware smoothness term popular in monocular depth estimation [godard2017unsupervised]. Specifically, we weigh depth gradients with the local intensity gradients in the RGB image, thereby encouraging smooth depth predictions while preserving discontinuities at actual object boundaries:
@@
Panoptic-edge-aware smoothness loss. In our multi-task setup, we leverage ground-truth panoptic segmentation labels to guide depth estimation. In depth maps, sharp discontinuities often coincide with object boundaries, which include both semantic class boundaries and instance-level boundaries within the same class. Previous semantic-aware smoothness losses [chen2019towards] only handle class transitions, e.g. between “road” and “car”, while several scenes also contain multiple adjacent objects of the same class, e.g. two adjacent “cars”, whose shared boundaries are equally informative about depth transitions. To overcome this limitation, we extend the semantic-edge formulation to a panoptic-edge one, ensuring that depth gradients respect both class and instance boundaries. In particular, we compute a binary mask based on the panoptic labels to determine where depth gradients should be penalized. More formally, let
@@
where $\mathcal{L}_{\text{seg}}$ denotes the full segmentation loss used in OneFormer [jain2023oneformer], which supervises our predicted semantics, $\mathcal{L}_{\text{cond}}$ supervises the CT embedding following CAFuser [cafuser], and $\mathcal{L}_{\text{depth}}$ supervises our predicted depth.
@@
Mask2Former [cheng2022masked] C Swin-T 49.35 39.38 48.84 46.48 45.39 45.1 31.29 58.23 78.34 58.01 46.89
MaskDINO [li2023mask] C Swin-T 51.85 42.73 54.05 46.20 46.23 48.54 38.64 57.29 80.51 60.28 49.44
OneFormer [jain2023oneformer] C Swin-T 57.55 47.83 58.33 53.68 53.43 53.77 43.54 63.69 80.93 66.97 55.21
HRFuser [broedermann2023hrfuser] CLRE HRFuser-T 44.60 40.01 47.03 43.59 42.69 40.62 28.33 55.23 78.41 54.34 43.90
MUSES [MUSES] CLRE 4xSwin-T 54.06 49.73 55.28 50.34 53.77 50.51 39.94 63.53 81.06 65.02 53.60
CAFuser-CAA [cafuser] CLRE Swin-T 59.93 56.24 61.16 56.41 59.38 57.88 48.03 67.64 81.80 71.61 59.38
CAFuser [cafuser] CLRE Swin-T 59.49 57.34 61.36 57.52 59.63 57.2 48.42 67.90 82.03 71.75 59.70

### v3

@@ Abstract
Index Terms:
Sensor fusion, semantic scene understanding, deep learning for visual perception
@@ I Introduction
Robust segmentation for automated driving systems demands reliable sensor fusion under any environmental condition. Conventional semantic perception pipelines that rely on a single sensor struggle in domains such as fog, rain, or low light, where the performance of individual sensors degrades. Prior work [4] has shown that conditioning cross-modal fusion on a global Condition Token (CT) that encodes the scene-wide environmental condition can effectively guide fusion by dynamically adapting sensor contributions. However, while this global approach models the overall scene condition, it overlooks the spatial variability of the effect of this condition on the sensor measurements, which largely depends on the local depth.
@@
In practice, depth is a key factor affecting the variable reliability of each sensor in each part of the scene. For instance, lidars provide accurate returns for nearby objects but degrade rapidly at longer distances, e.g. in fog or snowfall [15, 14]. Radars, on the other hand, operate well at long distances even in adverse weather but suffer from strong noise due to multi-path effects, whereas cameras are reliable only in well-lit parts of the scene, which are in turn related to the 3D structure of the scene. In other words, environmental conditions introduce spatial variations in the signal-to-noise ratio of each sensor, which call for a fine-grained local adaptivity of the fusion strategy. Consequently, incorporating explicit depth cues into the fusion mechanism for semantic perception could enable such spatial adaptivity to sensor reliability, which would in turn enhance segmentation robustness in challenging environments. This hypothesis is reinforced by several works on RGB-only semantics and depth multi-tasking [5, 26], which improve segmentation performance over semantics-only learning. Despite this promise for depth-informed fusion, most state-of-the-art multi-sensor segmentation methods [3, 29, 20, 18] treat lidar solely as an input modality, overlooking its potential for improving segmentation via multi-tasking by leveraging its pixel-aligned sparse depth measurements as ground truth for learning depth together with semantics.
@@
We fill in this gap and propose DGFusion, a depth-guided multimodal fusion network which upgrades condition-aware sensor fusion for semantic perception with explicit local depth guidance (cf. Fig. 1). Building on CAFuser [4], we adopt its shared backbone with modality-specific feature adapters and its global CT that encodes scene-wide environmental conditions. Beyond this, in DGFusion we introduce an auxiliary depth prediction head that leverages lidar’s pixel-aligned sparse depth measurements as noisy supervision to enrich local features with implicit depth cues to guide each local-window cross-attention fusion. This dual encoding of global environmental conditions and spatially-varying depth cues enables the network to make more precise decisions based on which sensor to rely more on for each feature. By incorporating depth cues, the fusion mechanism can adaptively adjust the weighting of each modality per image region, favoring the per-case more reliable sensor. Moreover, since these noisy depth measurements are available as one of our network’s inputs, i.e. the lidar, depth supervision comes at no extra cost compared to existing fusion pipelines. The synergy between the global condition and local depth representation is especially beneficial in adverse conditions, which have been found [1] to cause pronounced spatial variations in sensor reliability.
@@
Besides, training our model on adverse conditions poses additional challenges. For example, in fog or snowfall the lidar measurements, which provide pixel-aligned depth supervision, become inherently noisy [15, 14], making the direct integration of standard depth estimation approaches and losses in our multi-tasking model infeasible. In order to get a reliable training signal across all conditions, we carefully design our depth loss to be robust to sparse and noisy ground truth by combining an outlier-robust $L_{1}$ loss with strategies commonly used in monocular depth estimation. These strategies consist of an edge-aware smoothness loss to penalize depth gradients in uniform regions of the RGB image and a novel panoptic-edge-aware smoothness loss that better captures depth discontinuities between instances of the same semantic classes compared to previous semantic-edge-aware smoothness losses.
@@
Casting the training of DGFusion as a multi-tasking problem leverages the synergy between depth and semantics with minimal changes to the inference pipeline. Extensive experiments on MUSES [2] and DeLiVER [29] demonstrate that DGFusion sets the new state of the art in multi-sensor semantic perception across diverse conditions.
@@ II Related Work
Semantic perception encompasses tasks like semantic [10], panoptic [19], and instance [16] segmentation, which are pivotal for robotics applications such as automated driving [32]. Recent mask-based networks have advanced the state of the art [21], and unifying these tasks within a single model has proven beneficial [17]. In contrast to these RGB-only methods, our work focuses on multimodal semantic perception, leveraging sensor fusion to enhance robustness in challenging conditions where RGB data is unreliable.
@@
Sensor fusion is central in practical vision applications, as different sensors feature complementary characteristics. The availability of large multimodal datasets, from early ones like KITTI [12] to recent ones focusing on adverse conditions [29, 1, 2], has driven the development of fusion methods. Architectures have evolved from two-modality systems [7] to modular, attention-based networks that handle an arbitrary number of inputs [3, 29]. Recent approaches have integrated large-scale pre-trained models [20], explored modality-agnostic designs with shared backbones [31], and combined intra-modal with inter-modal attention [18]. The MUSES method [2] extends [9] to the multimodal setting, solving panoptic and semantic segmentation via local-window cross-attention. However, all these works fuse sensors uniformly, lacking explicit adaptation to the conditions that affect each image region non-uniformly, largely depending on depth.
@@
Condition-aware fusion utilizes information about the environmental conditions to guide perception. [7] addresses object detection by implicitly assigning larger weights to the modality that yields better detection features at late fusion. Knowledge distillation from vision-language models is used in [30] for learning modality-agnostic representations. The most closely related work to ours in this regard is [4], which guides cross-modal attention with a global camera-based CT that is learned contrastively based on training-time supervision with image-level verbal condition descriptions. In contrast, we aim at a more fine-grained, pixel-level adaptivity to the environmental condition, as its effects on sensor data vary non-uniformly with depth.
@@
Semantics and depth are strongly correlated tasks, and it is well-established that they benefit from being learned jointly in a multi-task setup [5, 26]; cf. [27] for a comprehensive overview of such semantics-depth multi-tasking methods. Prior work has leveraged this synergy, with semantics guiding monocular depth estimation [8, 33]. Similarly, depth supervision has benefited from edge-aware smoothness losses based on image gradients [13] and semantics-aware losses that use semantic annotations [22], which we extend to use more fine-grained panoptic annotations. In contrast to previous works, we explore the use of depth in multimodal settings as a readily available input signal, e.g. from sparse lidar measurements in typical outdoor multi-sensor suites, in order to benefit semantics. While the aforementioned multi-sensor semantic perception works utilize such depth information only as input to the fusion architecture, we instead leverage it as a reference output during training, reframing sensor fusion as a multi-task problem where depth completion serves as an auxiliary task to learn informative, localized features for semantic perception. The depth features learned from this auxiliary task are leveraged as a spatially varying conditioning in our attentive fusion together with global condition information. This combination allows DGFusion to adaptively weigh modalities not only per-input, but also per-region, based on how depth modulates the effect of the environmental condition on each sensor.
@@ III-A Architecture
Our goal is to achieve robust multimodal segmentation in all weather conditions. We build on CAFuser [4], which processes all modalities through a shared backbone with lightweight modality-specific feature adapters to align heterogeneous inputs into a common feature space, and uses global environmental condition encodings to guide the sensor fusion for the OneFormer [17] segmentation head. Following CAFuser, we project the secondary modalities (lidar, radar, and event camera) as 3-channel images onto the RGB camera plane and dilate them with a kernel of size $K_{\text{sensor}}$ to compensate for sparsity. We adopt CAFuser’s shared backbone with feature adapters and its global condition token (CT). In contrast to CAFuser, we treat lidar not only as an input modality but also as sparse, noisy ground-truth depth for multi-task learning, which poses additional challenges in adverse conditions where lidar measurements degrade. With our carefully designed robust losses (cf. Sec. III-B), this enables our network to learn depth-informed features that provide additional supervision for semantics and enable locally adaptive depth-guided fusion.
In Fig. 2, our complete DGFusion architecture is depicted. On the left, an arbitrary number of sensor modalities are processed by the shared backbone with modality-specific feature adapters, resulting in per-level feature maps. These features are then split into three branches: the depth estimation branch (top), the segmentation branch (center), and the condition representation branch adopted from CAFuser (bottom). The depth branch uses raw, undilated lidar ground truth to supervise depth predictions and induce depth-informed features via a lightweight fusion module, while the condition branch generates a global CT from the RGB input. Both branches provide extra cues to the central depth-guided fusion module.
@@
We use a Semantic FPN [19] decoder as our auxiliary depth head, which upsamples each depth feature $\mathbf{d}_{l}$ to 1/4 resolution and aggregates them:
@@
where Upsample consists of multiple convolutions and bilinear upsampling steps. This structure fuses coarse and fine information efficiently, enabling robust depth estimation. At inference, we drop the auxiliary depth head but keep the depth features $\mathbf{d}_{l}$ for guiding fusion in the semantic branch.
Condition representation branch. We adopt CAFuser’s [4] strategy to produce a global CT. We flatten the highest-level RGB features and process them with a lightweight Transformer composed of 2 encoder and 2 decoder layers:
@@
supervised during training with a verbo-visual contrastive loss [4, 17, 6] that leverages text prompts describing the environmental conditions in detail.
Depth-guided fusion. We use the depth features $\mathbf{d}_{l}$ to guide sensor fusion in our main semantic branch. Our Depth-Guided Fusion module utilizes parallel multi-window cross-attention fusion [3] to fuse in parallel each secondary modality via cross-attention with the primary camera modality within $K_{w}\times K_{w}$ local windows. This fusion design is efficient and scales well with the number of input modalities. For each level $l$ and secondary modality $m$, we partition RGB features $\mathbf{F}_{\text{rgb}}^{l}$, secondary features $\mathbf{F}_{\text{sec}}^{l,m}$, and depth features $\mathbf{d}_{l}$ into $I_{l}$ windows of size $K_{w}\times K_{w}$. For each window $i\in\{1,\ldots,I_{l}\}$, we extract a local Depth Token (DT):
@@ III-B Losses
Outlier-robust $L_{1}$ loss. Lidar returns degrade in adverse conditions, making the raw depth measurements noisy and unreliable. To address this, we compute an $L_{1}$ loss on logarithmic depth values. The $L_{1}$ loss is inherently more robust to outliers than $L_{2}$, and using the logarithmic scale better handles depth variations across orders of magnitude [26, 25, 24]. Let $\hat{D}_{p}$ be the predicted depth at pixel $p$ and $D_{p}$ the corresponding raw lidar reference depth. Since raw lidar depths are often severely corrupted in adverse conditions, directly supervising depth predictions using all available measurements can degrade learning (cf. Sec. IV-B). To address this, we restrict the loss computation to a more reliable subset of pixels, i.e. those pixels below the $\tau$-quantile of depth errors. We first define the absolute log-error at pixel $p$ as
@@
Edge-aware smoothness loss. To guide depth prediction in regions lacking reliable lidar supervision, we employ an edge-aware smoothness term popular in monocular depth estimation [13]. Specifically, we weigh depth gradients with the local intensity gradients in the RGB image, thereby encouraging smooth depth predictions while preserving discontinuities at actual object boundaries:
@@
Panoptic-edge-aware smoothness loss. In our multi-task setup, we leverage ground-truth panoptic segmentation labels to guide depth estimation. In depth maps, sharp discontinuities often coincide with object boundaries, which include both semantic class boundaries and instance-level boundaries within the same class. Previous semantic-aware smoothness losses [8] only handle class transitions, e.g. between “road” and “car”, while several scenes also contain multiple adjacent objects of the same class, e.g. two adjacent “cars”, whose shared boundaries are equally informative about depth transitions. To overcome this limitation, we extend the semantic-edge formulation to a panoptic-edge one, ensuring that depth gradients respect both class and instance boundaries. In particular, we compute a binary mask based on the panoptic labels to determine where depth gradients should be penalized. More formally, let
@@
where $\mathcal{L}_{\text{seg}}$ denotes the full segmentation loss used in OneFormer [17], which supervises our predicted semantics, $\mathcal{L}_{\text{cond}}$ supervises the CT embedding following CAFuser [4], and $\mathcal{L}_{\text{depth}}$ supervises our predicted depth.
@@
Mask2Former [9] C Swin-T 49.35 39.38 48.84 46.48 45.39 45.1 31.29 58.23 78.34 58.01 46.89
MaskDINO [21] C Swin-T 51.85 42.73 54.05 46.20 46.23 48.54 38.64 57.29 80.51 60.28 49.44
OneFormer [17] C Swin-T 57.55 47.83 58.33 53.68 53.43 53.77 43.54 63.69 80.93 66.97 55.21
HRFuser [3] CLRE HRFuser-T 44.60 40.01 47.03 43.59 42.69 40.62 28.33 55.23 78.41 54.34 43.90
MUSES [2] CLRE 4xSwin-T 54.06 49.73 55.28 50.34 53.77 50.51 39.94 63.53 81.06 65.02 53.60
CAFuser-CAA [4] CLRE Swin-T 59.93 56.24 61.16 56.41 59.38 57.88 48.03 67.64 81.80 71.61 59.38
CAFuser [4] CLRE Swin-T 59.49 57.34 61.36 57.52 59.63 57.2 48.42 67.90 82.03 71.75 59.70